OpenStudy (anonymous):

How can 1/3 x − 2 = 1/4 x + 11 be set up as a system of equations?

OpenStudy (anonymous):

Answer choices: 3y − x = −6 4y − x = 44 3y + x = −6 4y + x = 44 3y − 3x = −6 4y − 4x = 44 3y + 3x = −6 4y + 4x = 44

OpenStudy (midhun.madhu1987):

 Open the bracket, move all x terms to left side and the numbers to right side...